Write 2.85 as a percent.
Answer by stanbon(48526)

Percent means "over 100".
to change 2.85/1 to percent multiply numerator and denominator by 100to get:
(2.85/1)*(100/100) = 285/100 = 285 %
===================
Cheers,
Stan H.